pH measurements of a chemical solutions have mean 7.4 with standard deviation 0.01. Assuming all pH measurements of this solution have a nearly symmetric/bel curve distribution. Find the percent (%) of pH measurements reading below 739 OR above 7.4.please help me do this
Answer #1
for normal distribution z score =(X-μ)/σx
here mean=       μ= 7.4
std deviation   =σ= 0.0100

% of measurement below 7.39 or above 7.40 =1-P(7.39<X<7.40)=1-P(-1<Z<0)

=1-(0.5-0.1587)=0.6587 ~ 65.87%